OpenStudy (anonymous):

i love this website i will come back 2morrow its already 11:40 pm. gonna wake up 2morrow at 6:30 am. how many hours\minute do i have left to sleep?

OpenStudy (chrisplusian):

more than I ever get

hartnn (hartnn):

11.40-6.30 = 0.20+6.30 = 6.50 means 6 hours 50 minutes

OpenStudy (anonymous):

Anyway have a good sleep... :D

OpenStudy (anonymous):

oh my gosh i need 8 hour eye bugs time